TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-02-26 22:19:00
subject: 2\10 Pt-1 ISS On Orbit Status 10-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\10 ISS On-Orbit Status 10-02-2003
Part 1 of 2

ISS On-Orbit Status 2/10/03

All ISS systems continue to function nominally, except as noted 
previously or below. Day 79 in space for the Expedition 6 crew. Week 
11 for Increment 6.

Before breakfast, all crewmembers completed another session of the 
periodic Russian medical experiment protocols PZEh-MO-7 (calf volume 
measurement) and PZEh-MO-8 (body mass measurement). FE-1 Nikolai 
Budarin set up the MO-8 "scales" equipment and later broke it down and 
stowed it away. [Calf measurements (left leg only) are taken with the 
ISOG device, a custom-sewn fabric cuff that fits over the calf, using 
the knee and lower foot as fixed reference points, to provide a rough 
index of deconditioning in zero-G and effectiveness of 
countermeasures. For determining body ass in zero-G, where things are 
weightless (but not massless), the Russian IM "scales" measure the 
inertial forces that arise during the oscillatory motion of a mass 
driven by two helical metering springs with known spring constants. By 
measuring the time period of each oscillation of the unknown mass (the 
crewmember) and comparing it to the period of a known mass, the 
crewmember's mass is calculated by the computer and displayed].

Later in the day, Budarin started another round of the Russian MBI-8
"Profilaktika" (countermeasures) fitness test series, today on the 
VELO (stationary bike ergometer). [The test is identical to the 
Russian MO-5 assessment, but in addition to the nominal test 
procedure, it calls for the use of the TEEM-100M gas analyzer, 
measurement of the lactate level in the subject's blood with the 
AccuSport device, and subjective evaluation of physical exertion 
levels during the test. Results were entered on a log sheet. TEEM and 
ECG (electrocardiograph) data were transferred to Laptop 3 and 
prepared for later downlink via Regul-Packet comm.]

Budarin and CDR Ken Bowersox continued unloading and stowing cargo 
from Progress 10P while keeping track of items for the IMS (inventory 
management system) database.

A PCS HDD (portable computer system/hard disk drive) swap scheduled 
for Bowersox today was deferred to next week.

FE-2/SO Don Pettit worked on EXPRESS Rack 2 (ER2), setting the ARIS 
(active rack isolation system) up for standard control, by removing 
the ARIS ICE (isolation characterization experiment) Type 1553 cable 
and replacing it with a shorting plug.

Preparations for troubleshooting the deactivated MSG (microgravity 
science glovebox) continue on the ground, and a tagup/conference was 
conducted with Don Pettit at 11:00am EST. [The working group 
established by the ESA/Industrial Team and the Inquiry Board at 
ESA/ESTEC (Noordwijk/Holland) is narrowing down the possible causes of 
the failure. It already appears that the circuit breaker tripping was 
not a hard short-circuit but due to a non-nominal overload on the 
12Vdc line, which also affected the data transmission within the MSG 
core facility. Until the failure is completely isolated, the Glovebox 
rack will not be powered up, to avoid "waterfall"-type failures in its 
different subsystems. ESA inputs to MSFC POC (Payload Operations 
Center) for the physical troubleshooting were expected not earlier 
than today. Actual on-orbit troubleshooting is scheduled for Wednesday 
(2/12), after another teleconference with Pettit tomorrow.]

Nikolai Budarin took an EVARM (EVA radiation monitoring) badge 
reading, first unstowing and powering up the EVARM Reader, letting it 
warm up, then reading all badges and finally turning the Reader off 
again. [These badge readings help characterize the on-board radiation 
trends prior to and after an EVA.]

Nikolai performed the daily routine task of SOSh life support system 
(incl.  ASU toilet system) maintenance, Don prepared the daily IMS 
"delta" file for updating the inventory database, and Sox conducted 
the daily status check of Lab payloads.

Pettit continued the daily monitoring of CO2 partial pressure (ppCO2) 
levels in the SM and Lab (two readings) with the CO2 monitoring kit 
(CDMK).

Budarin completed his routine 24-hr. checkup and watering of the 
Russian BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

MCC-Houston continued IMV (intermodular ventilation) troubleshooting. 
[One of the crew's task list items in support of this activity is to 
check the Node "butterfly" plate (valve) positions. The ground is also 
unpowering and repowering the IMV fans at the Node aft port, Node aft 
starboard, and Lab aft port positions fans to capture and compare 
their power (current) signatures.]

Using the Kodak DSC760 or Nikon D1 digital camera, Budarin had 15 min.
scheduled for today's Uragan ("hurricane") Earth observations. 
[Nikolai was advised to take a series of images of the Genaldon River 
valley and the Northern slope of Kazbek-Dzhimarai mountain plexus in 
Russia,- former location of the Kolka Glacier. Land-based survey and 
aerial survey of the area are being conducted today as well. An 
Inter-Departmental Commission is currently evaluating the status of 
Kolka and Maili Glaciers and the condition of the debris/ice cone 
associated with these glaciers. Based on the results of this research, 
the Commission will develop recommendations on the 
continuation/improvement or possible cessation of activities that are 
currently being carried out in the area. This is why the ISS-based 
imagery of the area is such an important part of this effort.]

Coming up tomorrow: (a) ISS reboost by Progress 10P [for 21-22 minutes 
at 6:34am EST, using 430 kg propellants for ~6 m/sec delta-V. Altitude 
increase of 10.4 km is calculated to protect a possible Shuttle launch 
in mid-June (Shuttle-compatible altitude: 320 km).] (b) Transfer of 20 
liters water from the Lab condensate tank to a CWC (contingency water 
container).]

Today's targets for the CEO (crew earth observations program) were 
Riyadh, Saudi Arabia (the Saudi capital lied just right of track), 
Lisbon, Portugal (ISS had a nadir pass over the Portuguese capital, 
situated just inland on the north shore of the Tagus River estuary), 
Mexico City, Mexico (crew had a nice, midday pass at nadir over the 
Mexican capital city), Tucson, Arizona (weather may be marginal by the 
time of this pass, otherwise it will be a nadir view in midday sun of 
this desert city), and Albuquerque, New Mexico (fair weather should 
hold over this target. Crew had a nadir view of the city located 
mostly east of the Rio Grande River).

CEO images can be viewed at the website http://eol.jsc.nasa.gov

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.